# A Systematic Survey on Large Language Models for Algorithm Design

### v2

@@ 4. Search Methods
Incorporating LLMs within a search framework significantly enhances LLMs’ utility in algorithm design, making it the preferred option (Ma et al., 2024b; Liu et al., 2024d; Romera-Paredes et al., 2024). This section categorizes existing research according to the search methods employed, presents recent advancements, and discusses some of the limitations.
@@ 4.1.1. Beam Search
Beam search explores multiple promising paths in the search space. For instance, in LLM-based prompt engineering (Zhou et al., 2022; Pryzant et al., 2023; Ma et al., 2024d), increasing the beam size significantly enhances performance (Zhou et al., 2022). Beyond prompt engineering, beam search is utilized in diverse applications: Text2Motion (Lin et al., 2023) employs it for robotic task planning, SayCanPay (Hazra et al., 2024) for action sequence planning in robots, enhancing speech-processing tasks, and in a confidence-aware method for structured data generation.
@@ 4.2. Single-point-based Search
Single-point-based search methods iteratively refine a solution by leveraging neighborhood structures (Liu et al., 2024f) or specific search directions (Pryzant et al., 2023). While these methods can produce satisfactory results within a reasonable number of iterations, they may struggle with maintaining diversity and robustness in their search processes.
@@ 4.2.1. Hillclimb
Hillclimb search iteratively searches for better results, where the new one is generated based on the last one and the better one is survived. A basic version of Hillclimb is investigated by Zhang et al. (2024b) for algorithm design, where a heuristic is iteratively refined. The reasoning over the existing status is usually adopted to enhance each search step. For example, Li et al. (2024b) leverage LLMs to automatically design dense reward functions for RL agents in environments with sparse rewards, which iteratively refines the reward function based on feedback from the agent’s interactions with the environment and Yang et al. (2023) develop a multi-module framework with innovative feedback mechanisms, demonstrating through both LLM-based and expert evaluations that LLMs can effectively produce scientific hypotheses that are both novel and reflective of reality.
@@ 4.2.2. Neighborhood Search
Neighborhood search methods explore the solution space by iteratively modifying a current solution to find an improved one in a structured neighborhood. Notably, the LLM-GS framework (Liu et al., 2024f) has introduced a scheduled neighborhood search method that leverages two distinct strategies: 1) Programmatic Neighborhood Generation, which involves generating a neighborhood of programs by selecting a node from the abstract syntax tree of a given program and substituting it with a subtree that is randomly generated following the production rules and sampling strategies of a domain-specific language; and 2) Latent Space Representation, where the neighborhood is defined in a latent space by training a variational autoencoder on a corpus of randomly generated DSL programs, thereby creating a more abstract and potentially more informative neighborhood structure. Additionally,  Jin et al. (2023) contribute to this line of research by defining neighborhood algorithms based on a minimum cosine similarity threshold of 60% between the embeddings of a query vector and code snippets, enhancing the search process by ensuring that algorithmic variations are semantically coherent and aligned with the underlying problem structure.
@@ 4.2.3. Gradient-based Search
Gradient-based search utilizes a (pseudo) gradient direction to generate new designs in each iteration. Unlike searches in a continuous space with a differentiable objective, calculating the gradient in the algorithm design space poses significant challenges. Consequently, a pseudo descent direction is often employed. Pryzant et al. (2023) introduce automatic prompt optimization, which improves prompts automatically by adjusting them based on natural language “gradients” derived from training data. Similarly, Tang et al. (2024c) introduce GPO, which draws parallels with gradient-based model optimizers for prompt optimization. Moreover, Nie et al. (2024) explore the use of LLMs as interactive optimizers for solving maximization problems in a text space through natural language and numerical feedback. For an accurate gradient, Guo et al. ([n. d.]) perform the gradient-based search in a mapped continuous space and proposes a collaborative optimization strategy that combines a gradient-based optimizer with an LLM for tackling complex non-convex optimization challenges. These works employ either pseudo gradients or gradients over mapped text space; however, a systematic study of gradients over general algorithm space is anticipated.
@@ 4.2.4. Reinforcement Learning
Reinforcement learning learns to make decisions by interacting with an environment to maximize rewards. GPTswarm (Zhuge et al., 2024) and Duan et al. (2023) leverage LLMs and RL techniques to optimize code, reducing complexity and enhancing efficiency. Similarly, Liu et al. (2024f) present a novel LLM-guided search framework (LLM-GS) aimed at addressing the sample inefficiency in state-of-the-art programmatic RL methods by utilizing the programming expertise of LLMs to boost search efficiency. Additionally, LLMs have been utilized in the design of RL systems, particularly in the creation of reward-shaping functions. Moreover, Bhambri et al. (2024) propose a framework called MEDIC to tackle the issue of sample inefficiency. They use LLMs to generate a guide policy for constructing reward-shaping functions.
@@ 4.3. Population-based Search
Population-based evolutionary search is the main tool investigated in LLM4AD papers due to its effectiveness and robustness in complex search space (O’Reilly and Hemberg, 2024; Wu et al., 2024a). The majority of these works use a simple genetic algorithm with greedy population management Liu et al. (2024d). Some of them investigate advanced population management including multi-island Romera-Paredes et al. (2024), quality-diversity Hu et al. (2024), and dominance-dissimilarity measurement Yao et al. (2024b). According to the number of objectives, they can be roughly categorized into single-objective search and multi-objective search.
@@ 4.3.1. Single-objective Evolutionary Search
On algorithm design tasks involving text and code search, the complicated search space poses challenges for diversity control and population management. The majority of workers adopt a greedy way (Liu et al., 2024d; Chen et al., 2024a; van Stein and Bäck, 2024; Li et al., 2024b; Narin, 2024; Guo et al., 2024b; Yao et al., 2024a). In these works, a population of individuals is maintained and only the ones with better fitness will survive. Romera-Paredes et al. (2024) adopt a multi-island evolutionary algorithm to explore a diverse population with a dynamically adjusted population size. Moreover, Wong et al. (2024a) use LLMs to create digital artifacts for creative and engineering applications, utilizing a multi-factorial evolutionary algorithm to drive an LLM. Quality diversity methods (Pugh et al., 2016) have also been investigated for this purpose, Lehman et al. (2023) and Hu et al. (2024) utilize the MAP-Elites algorithm with a grid of niches to generate diverse and high-quality solutions.
@@ 4.4. Uncertainty-guided Search
Uncertainty-guided search harnesses the strengths of Bayesian Optimal Experiment Design (BOED) or Bayesian Optimization (BO) frameworks within the search process incorporating LLMs. It first establishes a prior that describes the initial belief about the parameters of interest and then employs uncertainty-driven strategies over the derived posterior to iteratively update the belief, ultimately enabling efficient decision-making during the search process in a sample-efficient manner.
This search paradigm with uncertainty has demonstrated its superiority in some scenarios. For example, Handa et al. (2024) introduce OPEN using LLMs to extract environment-relevant features and translate abstract feature queries into natural language questions to interact with users through conversations. Meanwhile, BOED is employed to identify the most informative query for efficiently learning preference information. To improve the multi-turn and decision-theoretic reasoning capabilities over a set of arbitrary items, Austin et al. (2024) formulate the natural language-based preference elicitation within a BO framework. The proposed PEBOL algorithm establishes a probabilistic prior over user preferences in the form of Beta distributions. In the domain of LLM decoding, Grosse et al. (2024) propose a probabilistic model named ULTS. This method frames the decoding process as a tree-search problem and aims to find the optimal path in the LLM search tree by sequential decision-making under uncertainty.
@@ 5. Prompt Strategies
(a) Pre-trained and domain LLMs
@@
(c) LLMs
[…]
Figure 8. LLM types and promote engineering methods
Fig. 8(a) depicts the percentage of domain or pre-trained LLMs used in the literature. Among them, over 80% chose to use the pre-trained model without any specific finetuning, and about 10% fine-tuned the pre-trained model on domain datasets (Cummins et al., 2024) in which 4.4% are trained from scratch. Fig. 8(c) illustrates the top-used LLMs. LLM4AD papers show a strong preference for GPT models. GPT-4 and GPT-3.5 are the most-used LLMs, which in total take around 50%. Llama-2 is the most used open-source LLM. Once we have the pre-trained LLMs, prompt engineering is significant for effectively integrating LLMs in algorithm design. We introduce the main prompt engineering methods including zero-shot, few-shot, chain-of-thought, self-consistency, and reflection (Sahoo et al., 2024) used in LLM4AD papers. Fig. 8(b) shows the number of papers involving different prompt engineering techniques.
@@ 5.1. Zero-Shot
Zero-Shot (ZS) prompting enables a model to comprehend and perform tasks without prior specific training on those tasks. In algorithm design, zero-shot prompting allows direct requests to the LLM for solutions or responses, examples including conversational replies (Sekulić et al., 2024), assertion design (Miftah et al., 2024), RL guidance (Zhang et al., 2024e), and molecule generation (Ye, 2024). The model leverages its pre-trained knowledge to generate these responses. Although this method is advantageous for swiftly producing solutions based on general knowledge, it may not consistently deliver tailored responses to the nuanced demands of complex algorithmic challenges (Zhang et al., 2024b).
@@ 5.2. Few-Shot
Few-Shot (FS) prompting involves providing the model with a handful of examples to illustrate the task before it attempts to solve a similar problem. This approach enhances the model’s understanding of the context and its ability to tailor responses to specific tasks. In algorithm design, this method proves particularly effective by presenting the model with examples of algorithms (Liu et al., 2023b), solutions (Yang et al., 2024), prompts (Wang et al., 2023b; Tang et al., 2024c), and codes (Ma et al., 2024a; Morris et al., 2024). The examples used in the prompt may be manually-designed seeds (Romera-Paredes et al., 2024; Hu et al., 2024) or derived from LLM-generated designs (Liu et al., 2024d; Wong et al., 2024a; Ma et al., 2024c; Morris et al., 2024; Wang et al., 2023b; Tang et al., 2024c; Custode et al., 2024). These samples are typically sorted to suggest a preference to enhance performance (Yang et al., 2024; Romera-Paredes et al., 2024).
@@ 5.3. Chain-of-Thought
Chain-of-Thought (CoT) prompting encourages the model to articulate intermediate steps or reasoning paths that lead to the final answer. This technique is particularly valuable in algorithm design, where understanding the step-by-step process or engaging in instructed reasoning over existing designs helps prevent outlier results and fosters effective algorithm development. For instance, Liu et al. (2024d) introduce several prompting strategies, one of which directs the LLM to reason over existing heuristics, summarize the general pipeline, and design a new one based on this reasoning. Custode et al. (2024) prompt the LLM to evaluate whether the current step size is sufficient for convergence and to suggest adjustments based on its reasoning. In addition, multiple LLM-based agents are employed within a CoT framework by Sun et al. (2024b) to enhance heuristic design for a boolean satisfiability problem solver.
@@ 5.4. Self-consistency
Self-consistency involves generating multiple answers or solutions to the same prompt and then synthesizing them to improve accuracy and reliability. In algorithm design, this could mean prompting the model multiple times for different approaches to solve a problem and then comparing these solutions to identify the most efficient or robust algorithm. This approach leverages the model’s ability to generate diverse solutions and builds a more comprehensive understanding of possible strategies. For example, Guan et al. (2023) set the number of responses as 10 and uses the majority vote to get the predicted label. As each response may provide different keywords or regular expressions, it takes the union of the keywords or regular expressions to create a candidate set. Moreover, Lehman et al. (2024) sample multiple revised codes from the same selected code and updates the population accordingly.
@@ 5.5. Reflection
Reflection in prompt engineering involves asking the model to critique or evaluate its own responses or solutions. After generating an algorithm/solution, the model can be prompted to reflect on the efficiency, potential flaws, or improvements. Ma et al. (2024d) investigate both implicit and explicit reflection in LLM-driven prompt optimization, which lets LLM analyze the errors and generate a reflection or feedback regarding the current prompt. An additional study by Ye et al. (2024b) incorporates short-term and long-term reflections in heuristic design, while Zhang et al. (2024c) focus on self-evaluation and reasoning over results from network detection. Furthermore, both Ma et al. (2024b) and Narin (2024) adopt LLM for RL reward function design with a reward reflection that monitors the scalar values of all reward components and the task fitness function at various policy checkpoints during training.
@@ 6.1.1. Combinatorial Optimization
In the domain of Combinatorial Optimization (CO), automatic algorithm heuristics design has been a significant area of interest for a long time. The Traveling Salesman Problem (TSP) stands out as one of the most renowned CO problems, involving the quest for the shortest route to visit all specified locations exactly once and return to the starting point. Some recent work leverages LLMs to evolve algorithms within Evolutionary Computation (EC) framework, such as AEL (Liu et al., 2023b), ReEvo (Ye et al., 2024a) and EoH (Liu et al., 2024d). Differently, OPRO (Yang et al., 2024) employs LLMs as optimizers with a proposed meta-prompt, in which the solution-score pairs with task descriptions are added in each optimization step. Additionally, LMEA (Liu et al., 2024b) investigates the utilization of LLMs as evolutionary combinatorial optimizers for generating offspring solutions, wherein a self-adaptation mechanism is introduced to balance exploration and exploitation. The Capacitated Vehicle Routing Problem (CVRP) extends the TSP by introducing constraints related to vehicle capacity. To address this challenge, MLLM (Huang et al., 2024c) devises a multi-modal LLM-based framework with textual and visual inputs to enhance optimization performance.
In addition to routing problems, other combinatorial optimization problems that have been investigated include cap set (Romera-Paredes et al., 2024), bin packing (Liu et al., 2024d), flow job shop scheduling (Liu et al., 2024d) and social networks problems (Sartori et al., 2024).

### v3

@@ 4. Search Methods
Incorporating LLMs within a search framework significantly enhances LLMs’ utility in algorithm design, making it the preferred option (Ma et al., 2024b; Liu et al., 2024e; Romera-Paredes et al., 2024). This section categorizes existing research according to the search methods employed, presents recent advancements, and discusses some of the limitations.
@@ 4.1.1. Beam Search
Beam search explores multiple promising paths within a search space. For example, in the context of LLM-based prompt engineering (Zhou et al., 2022; Pryzant et al., 2023; Ma et al., 2024d), increasing the beam size has been shown to significantly improve performance. Beyond prompt engineering. Beam search is also employed in a variety of other applications. For instance, Text2Motion (Lin et al., 2023) uses beam search for robotic task planning. Similarly, SayCanPay (Hazra et al., 2024) applies it to plan action sequences in robotics.
@@ 4.2. Single-point-based Search
Single-point-based search methods iteratively refine a solution by leveraging neighborhood structures (Liu et al., 2024g) or specific search directions (Pryzant et al., 2023). While these methods can produce satisfactory results within a reasonable number of iterations, they may struggle with maintaining diversity and robustness in their search processes.
@@ 4.2.1. Hillclimb
Hillclimb search iteratively searches for better results, where the new one is generated based on the last one and the better one is survived. A basic version of Hillclimb is investigated by Zhang et al. (2024b) for algorithm design, where a heuristic is iteratively refined. The reasoning over the existing status is usually adopted to enhance each search step. For example, Li et al. (2024b) leverage LLMs to automatically design dense reward functions for RL agents in environments with sparse rewards, which iteratively refines the reward function based on feedback from the agent’s interactions with the environment. Yang et al. (2023) develop a multi-module framework with innovative feedback mechanisms, demonstrating through both LLM-based and expert evaluations that LLMs can effectively produce scientific hypotheses that are both novel and reflective of reality.
@@ 4.2.2. Neighborhood Search
Neighborhood search methods explore the solution space by iteratively modifying a current solution to find an improved one in a structured neighborhood. Notably, the LLM-GS framework (Liu et al., 2024g) has introduced a scheduled neighborhood search method that leverages two distinct strategies: 1) Programmatic Neighborhood Generation, which involves generating a neighborhood of programs by selecting a node from the abstract syntax tree of a given program and substituting it with a subtree that is randomly generated following the production rules and sampling strategies of a domain-specific language; and 2) Latent Space Representation, where the neighborhood is defined in a latent space by training a variational autoencoder on a corpus of randomly generated DSL programs, thereby creating a more abstract and potentially more informative neighborhood structure. Additionally,  Jin et al. (2023) contribute to this line of research by defining neighborhood algorithms based on a minimum cosine similarity threshold of 60% between the embeddings of a query vector and code snippets, enhancing the search process by ensuring that algorithmic variations are semantically coherent and aligned with the underlying problem structure.
@@ 4.2.3. Gradient-based Search
Gradient-based search utilizes a (pseudo) gradient direction to generate new designs in each iteration. Unlike searches in a continuous space with a differentiable objective, calculating the gradient in the algorithm design space poses significant challenges. Consequently, a pseudo descent direction is often employed. Pryzant et al. (2023) introduce automatic prompt optimization, which improves prompts automatically by adjusting them based on natural language “gradients” derived from training data. Similarly, Tang et al. (2024c) introduce GPO, which draws parallels with gradient-based model optimizers for prompt optimization. Moreover, Nie et al. (2024) explore the use of LLMs as interactive optimizers for solving maximization problems in a text space through natural language and numerical feedback. For an accurate gradient, Guo et al. ([n. d.]) perform the gradient-based search in a mapped continuous space and propose a collaborative optimization strategy that combines a gradient-based optimizer with an LLM for tackling complex non-convex optimization challenges. These works employ either pseudo gradients or gradients over mapped text space, however, a systematic study of gradients over general algorithm space is anticipated.
@@ 4.2.4. Reinforcement Learning
Reinforcement learning learns to make decisions by interacting with an environment to maximize rewards. Zhuge et al. (2024) and Duan et al. (2023) leverage LLMs and RL techniques to optimize code, reducing complexity and enhancing efficiency. Similarly, Liu et al. (2024g) present a novel LLM-guided search framework called LLM-GS, which aims at addressing the sample inefficiency in state-of-the-art programmatic RL methods by utilizing the programming expertise of LLMs to boost search efficiency. Additionally, LLMs have been utilized in the design of RL systems, particularly in the creation of reward-shaping functions. Moreover, Bhambri et al. (2024) use LLMs to generate a guide policy for constructing reward-shaping functions to tackle the issue of sample inefficiency.
@@ 4.3. Population-based Search
Population-based evolutionary search is the main tool investigated in LLM4AD papers due to its effectiveness and robustness in complex search space (O’Reilly and Hemberg, 2024; Wu et al., 2024a). The majority of these works use a simple genetic algorithm with greedy population management (Liu et al., 2024e). Some of them investigate advanced population management including multi-island (Romera-Paredes et al., 2024), quality-diversity (Hu et al., 2024), and dominance-dissimilarity measurement (Yao et al., 2024b).
@@ 4.3.1. Single-objective Evolutionary Search
On algorithm design tasks involving text and code search, the complicated search space poses challenges for diversity control and population management. The majority of works adopt a greedy way (Liu et al., 2024e; Chen et al., 2024a; van Stein and Bäck, 2024; Li et al., 2024b; Narin, 2024; Guo et al., 2024b; Yao et al., 2024a; Hemberg et al., 2024). In these works, a population of individuals is maintained and only the ones with better fitness will survive. Romera-Paredes et al. (2024) adopt a multi-island evolutionary algorithm to explore a diverse population with a dynamically adjusted population size. Moreover, Wong et al. (2024a) use LLMs to create digital artifacts for creative and engineering applications, utilizing a multi-factorial evolutionary algorithm to drive an LLM. Quality diversity methods (Pugh et al., 2016) have also been investigated for this purpose, Lehman et al. (2023) and Hu et al. (2024) utilize the MAP-Elites algorithm with niches to generate diverse and high-quality solutions.
@@ 4.4. Uncertainty-guided Search
Uncertainty-guided search combines Bayesian Optimal Experiment Design (BOED) or Bayesian Optimization (BO) with Large Language Models (LLMs). It starts with a prior based on initial parameter beliefs, then uses uncertainty-driven strategies to iteratively refine the beliefs, enhancing decision-making efficiency in a sample-efficient way. This method has shown effectiveness in various applications. For example,  Handa et al. (2024) develop OPEN, which uses LLMs to extract environmental features and translate feature queries into conversational natural language questions, employing BOED to select the most informative queries for learning user preferences efficiently. Austin et al. (2024) introduce the PEBOL algorithm within a BO framework to improve multi-turn and decision-theoretic reasoning, using Beta distributions to model user preferences probabilistically. In LLM decoding, Grosse et al. (2024) propose ULTS, which treats decoding as a tree-search problem and optimizes path selection through sequential decision-making under uncertainty.
@@ 5. Prompt Strategies
Prompt strategies are vital for effectively utilizing LLMs, particularly in algorithm design tasks that require instructions on reasoning and reflection on the target tasks. We will begin by providing an overview of LLMs and the prompt strategies employed in existing works, followed by a review to works using each strategy.
Fig. 9(a) depicts the percentage of domain or pre-trained LLMs used in the literature. Among them, over 80% choose to use the pre-trained model without any specific finetuning, and about 10% fine-tuned the pre-trained model on domain datasets (Cummins et al., 2024) in which 4.4% are trained from scratch. LLM4AD papers show a strong preference for GPT models. GPT-4 and GPT-3.5 are the most commonly utilized LLMs, collectively accounting for approximately 50%. Llama-2 is the most used open-source LLM. Once we have the pre-trained LLMs, prompt engineering is significant for effectively integrating LLMs into algorithm design. LLM4AD papers involve many prompt engineering methods (as illustrated in Fig. 9(b)) including Zero-Shot (ZS), Few-Shot (FS), Chain-of-Thought (CoT), Self-Consistency (SC), and Reflection (RF) (Sahoo et al., 2024).
(a) LLMs
@@
Figure 9. LLM types and prompt engineering methods
@@ 5.1. Zero-Shot
Zero-Shot (ZS) prompting enables a model to comprehend and perform tasks without prior specific training on those tasks. In algorithm design, zero-shot prompting allows direct requests to the LLM for solutions or responses, examples including conversational replies (Sekulić et al., 2024), assertion design (Miftah et al., 2024), RL guidance (Zhang et al., 2024f), and molecule generation (Ye, 2024). The model leverages its pre-trained knowledge to generate these responses. Although this method is advantageous for swiftly producing solutions based on general knowledge, it may not consistently deliver tailored responses to the nuanced demands of complex algorithmic challenges (Zhang et al., 2024b).
@@ 5.2. Few-Shot
Few-Shot (FS) prompting involves providing the model with a handful of examples to illustrate the task before it attempts to solve a similar problem. This approach enhances the model’s understanding of the context and its ability to tailor responses to specific tasks. In algorithm design, this method proves particularly effective by presenting the model with examples of algorithms (Liu et al., 2023b), solutions (Yang et al., 2024), prompts (Wang et al., 2023b; Tang et al., 2024c), and codes (Ma et al., 2024a; Morris et al., 2024). The examples used in the prompt may be manually-designed seeds (Romera-Paredes et al., 2024; Hu et al., 2024) or derived from LLM-generated designs (Liu et al., 2024e; Wong et al., 2024a; Ma et al., 2024c; Morris et al., 2024; Wang et al., 2023b; Tang et al., 2024c; Custode et al., 2024). These samples are typically sorted to suggest a preference to enhance performance (Yang et al., 2024; Romera-Paredes et al., 2024).
@@ 5.3. Chain-of-Thought
Chain-of-Thought (CoT) prompting encourages the model to articulate intermediate steps or reasoning paths that lead to the final answer. This technique is particularly valuable in algorithm design, where understanding the step-by-step process or engaging in instructed reasoning over existing designs helps prevent outlier results and fosters effective algorithm development. For instance, Liu et al. (2024e) introduce several prompting strategies, one of which directs the LLM to reason over existing heuristics, summarize the general pipeline, and design a new one based on this reasoning. Custode et al. (2024) prompt the LLM to evaluate whether the current step size is sufficient for convergence and to suggest adjustments based on its reasoning. In addition, multiple LLM-based agents are employed within a CoT framework by Sun et al. (2024b) to enhance heuristic design for a boolean satisfiability problem solver.
@@ 5.4. Self-consistency
Self-Consistency (SC) involves generating multiple answers or solutions to the same prompt and then synthesizing them to improve accuracy and reliability. In algorithm design, this could mean prompting the model multiple times for different approaches to solve a problem and then comparing these solutions to identify the most efficient or robust algorithm. This approach leverages the model’s ability to generate diverse solutions and builds a more comprehensive understanding of possible strategies. For example, Guan et al. (2023) set the number of responses as 10 and uses the majority vote to get the predicted label. As each response may provide different keywords or regular expressions, it takes the union of the keywords or regular expressions to create a candidate set. Moreover, Lehman et al. (2024) sample multiple revised codes from the same selected code and updates the population accordingly.
@@ 5.5. Reflection
Reflection (RF) in prompt engineering involves asking the model to critique or evaluate its own responses or solutions. After generating an algorithm/solution, the model can be prompted to reflect on the efficiency, potential flaws, or improvements. Ma et al. (2024d) investigate both implicit and explicit reflection in LLM-driven prompt optimization, which lets LLM analyze the errors and generate a reflection or feedback regarding the current prompt. An additional study by Ye et al. (2024b) incorporates short-term and long-term reflections in heuristic design, while Zhang et al. (2024c) focus on self-evaluation and reasoning over results from network detection. Furthermore, both Ma et al. (2024b) and Narin (2024) adopt LLM for RL reward function design with a reward reflection that monitors the scalar values of all reward components and the task fitness function at various policy checkpoints during training.
@@ 6.1.1. Combinatorial Optimization
In the domain of Combinatorial Optimization (CO), automated algorithm heuristics design has been a significant area of interest for a long time. The Traveling Salesman Problem (TSP) stands out as one of the most renowned CO problems, involving the quest for the shortest route to visit all specified locations exactly once and return to the starting point. Some recent work leverages LLMs to evolve algorithms within Evolutionary Computation (EC) framework, such as AEL (Liu et al., 2023b), ReEvo (Ye et al., 2024a) and EoH (Liu et al., 2024e). Differently, OPRO (Yang et al., 2024) employs LLMs as optimizers with a proposed meta-prompt, in which the solution-score pairs with task descriptions are added in each optimization step. Additionally, LMEA (Liu et al., 2024b) investigates the utilization of LLMs as evolutionary combinatorial optimizers for generating offspring solutions, wherein a self-adaptation mechanism is introduced to balance exploration and exploitation. The Capacitated Vehicle Routing Problem (CVRP) extends the TSP by introducing constraints related to vehicle capacity. To address this challenge, MLLM (Huang et al., 2024c) devises a multi-modal LLM-based framework with textual and visual inputs to enhance optimization performance. In addition to routing problems, other combinatorial optimization problems that have also been investigated include cap set (Romera-Paredes et al., 2024), bin packing (Liu et al., 2024e), flow job shop scheduling (Liu et al., 2024e) and social networks problems (Sartori et al., 2024).